OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2 A. x = (x – 6) + 2 B. (y + 2) + y = 6 C. x + (y + 2) = 6 D. x + y = y + 2

OpenStudy (texaschic101):

x + y = 6 x = y + 2 now sub y + 2 in for x in the 1st equation (y + 2) + y = 6

OpenStudy (anonymous):

Thank you so much